Suppose you are one of 17p eople at a dinner party. Find the probability that at least one of the other guests has the same birthday as you and that some pair of guests share the same birthday. Assume there are 365 days in the year.

The probability that at least one other guest shares your birthday is approximately _____(Round to four decimal places as needed.)

The probability that some pair of guests share the same birthday is approximately _____(Round to four decimal places as needed.)
Answer #1

1) The probability that at least one other guest shares your birthday is approximately =1-P(none share the same birthday)

=1-(364/365)16 =  0.0429

2)probability that some pair of guests share the same birthday is approximately =1-P(nobody share the birthday)

=1-365C17/36517 =0.3150